OpenStudy (anonymous):

find derivative of (cscx)/(3sinx) I got (3cot+3)/(3sinx)^2 but I need verification.

OpenStudy (anonymous):

\[(-2/3) \cot(x) \csc^2(x)\] is what I get. Factor out constants: 1/3 (d/dx(csc^2(x))) Use the chain rule, d/dx(csc^2(x)) = ( du^2)/( du) ( du)/( dx), where u = csc(x) and ( du^2)/( du) = 2 u: 1/3 (2 csc(x) (d/dx(csc(x)))) The derivative of csc(x) is -cot(x) csc(x): -2/3 csc(x) (-cot(x) csc(x))

OpenStudy (anonymous):

we haven't learned chain rule yet, is there ayway to do it without chain rule?

OpenStudy (anonymous):

Have you learnt the quotient rule?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sin x cosec x =1 cosec x cos x = cot x Did you use these?

OpenStudy (anonymous):

Try taking a factor of (1/3) out of the equation before you use the quotient rule also, this may help.

OpenStudy (anonymous):

i dotn understand how you got that. this is what I did, I took out the constant 1/3, so now I have cscx/sin....so (sin)(-cscxcotx)-(cscx)(cosx)/(sinx)^2

OpenStudy (anonymous):

yep, you can simplify what you have, remember cosec x is just 1/sin x, so it will simplify to -cot x on the LHS of the subtraction sign...

OpenStudy (anonymous):

yeah I understand that part, now I have (-cotx-cscxcosx)/(sin)^2....so what else would i simplify

OpenStudy (anonymous):

since cosec is 1/sin, you have cos/sin which is cot, now you have cot cot which is cot^2

OpenStudy (anonymous):

the original answer is just rearranged slightly, you're good

OpenStudy (anonymous):

do you mean -cot^2?